# Emergent broadband polarization entanglement from electronic and phononic Stokes–anti-Stokes indistinguishability

### v2

@@ Abstract
With the values of $A^{\eta}_{ijkl}$ in hand, one can predict what is the entanglement in the SaS scattered state $|\Psi_{SaS}\rangle(\theta)$ of Eq. (3) for given values of $\omega_{S}$ and $\omega_{aS}$. We use, as an entanglement measure, the entropy of entanglement $E=-\text{Tr}_{i}(\rho_{i}\text{log}_{2}\rho_{i})$ of either subsystem $i=\{S,aS\}$ with reduced state $\rho_{i}$ [17], which, since our global state is pure, is zero for separable states and unity for maximally entangled states.
@@
Figure 2: Entanglement map for the SaS states in (a) the present experiment and (b) a generic centrosimetric cubic system. In (a), the laser bandwidth $W$ is fixed at $W=264\text{ cm}^{-1}$, the value of our experiment, as localized in (b) by a horizontal line. The black vertical line is the resonant Raman shift $\omega_{ph}/2\pi$, and the red regions mark the entanglement maxima ($E>0.999$). As tensor components in Eq. (5) we used the values from table 1. The star localizes the experiment of Ref. [12]. In blue shades, we plot contour curves for the Gisin parameter $F$. The hatched regions, within one laser bandwidth FWHM, indicates a high production of independent SaS photon pairs.
@@
In Fig. 2 we show a contour plot of $E(\omega_{S},\omega_{aS};\theta,W)$, under the condition that $\omega_{aS}=2\omega_{c}-\omega_{S}$, that is, a symmetric SaS Raman shift. On the horizontal axis we vary the Raman shift, going through the resonance in diamond at $\omega_{ph}/2\pi=1332$ cm-1, identified by a black vertical line, and on the vertical axis we plot (a) the angle $\theta$ between the laser linear polarization and the crystallographic axis, going from $\theta=0$ to $45^{\circ}$ (due to the symmetry of the crystal, all other $\theta$ values will be related to this range), and (b) the laser bandwidth $W$, going from zero to $W=190\gamma$, and where the value for our experiment $W=24\gamma$ ($W=264\text{ cm}^{-1}$ or FWHM 70 cm-1) is indicated by a black horizontal line. The thick red regions indicate the parameters where $E\rightarrow 1$, showing maximum entanglement.
@@
In Fig. 2 there is a hatched region indicating where the uncorrelated SaS pair production is high (low $g^{(2)}$ region in Fig. 1 (c,d)), corresponding to $2W$, around 1.2×1.2$\times$1.2 × FWHM. In this region, the scattered state is not properly represented by Eq. (2), but it needs to be complemented by other SaS scattering events, which involves the scattering of real phonons, that happens within a laser bandwidth around the Raman resonance [19]. Outside this region, though, state (2) is a good representation of the scattered SaS state, and our entanglement measure is representative.
@@
To complete the analysis, we have also drawn level curves for the violation of Bell-type CHSH inequalities as a function of the light-matter parameters. These curves correspond to the so-called Gisin parameter $F$ [20, 21] which, for pure states, reads $F=2(1-\mathcal{P})^{1/2}$, where $\mathcal{P}=1-\text{Tr}_{i}(\rho_{i})^{2}$ is the linear entropy of either subsystem $i$. For separable states $F=2$, which means that no separable state can go above the classical upper bound in the CHSH inequality, equal to 2. It reaches its maximum at $F=2\sqrt{2}\approx 2.83$ for maximally entangled states. We plot the $F$ parameter in Fig. 2 as blue shaded contour curves, and its smallest values within a white contour. Note that the maximal violation coincides with the maximally entangled states near 1200 cm-1 and 1400 cm-1. Since Bell analysis is state dependent, and the generated $|\Psi_{SaS}\rangle$ state depends on the values of the light-matter parameters, the appropriate Bell angles to reach the maximum violation change for different regions of the map. As an example, the violation in the maximally entangled states near 1200 cm-1 is obtained for linear polarization angles $(0,\pi/4)$ for one of the photons and $(\pi/8,3\pi/8)$ for the other, where the angles are in relation to $V$ polarization. Note that the quality of the entanglement, characterized by both the Gisin parameter $F$ and the entanglement entropy $E$, are independent of the rate of pair generation, so no matter how many SaS photon pairs are generated, our analysis yields how much entanglement will be available with each pair.
@@
Finally, we can localize the result of Ref. [12] in Fig. 2, represented by a white star. In the paper, it was found a state close to a Bell state at a symmetric Raman shift of 900 cm-1 with $\theta=0^{\circ}$, with amplitude ratio of $|HH\rangle$ to $|VV\rangle$ at $\sqrt{0.28/0.72}\approx 0.62$. Calculating the ratio with our state (3) yields 0.49, and the discrepancy comes from our aS filter cutting some of the Raman contribution in HH at 900 cm-1 (left edge of Fig. 1 (a)), which does not happen in Ref. [12]. This region can be seen in the bottom left of Fig. 2 (a) and (b), where and $E\approx 0.7$ and $F\approx 2.5$.
@@
To conclude, this work presents a fully quantum theory that describes the broadband scattered state in the SaS process, including photon pair polarization. It incorporates experimental conditions like the laser bandwidth and crystal orientation, to allow for the calculation of any sought physical quantity (e.g. $g^{(2)}(0)$ and entanglement). The efficiency of the process can be controlled by how close the system is to the resonance with a Raman active phonon, and it is a matter of state engineering to choose the appropriate laser–crystal angle and SaS frequency in order to obtain a good balance between a good amount of entanglement and a sufficient pair production rate. This theory has been elusive to find despite a decade-long effort [2, 10, 15, 22, 23, 12]. The result is explored here for diamond, but it should be general for other centrosymmetric media, including silicon, whose electronic band gap (1.12 eV) is one order of magnitude higher than the phonon energy ($<0.1$ eV), respecting the assumptions of our model. This should considerably bolster the engineering of SaS photon pair generation.
@@
For nonzero $\chi^{(2)}$ materials, spontaneous parametric down-conversion (SPDC) enables broadband pair generation as well, as it has been recently demonstrated [25]. However, in SPDC the generation of a polarization entangled photon pair is due to the decay of a single highly energetic photon [5], demanding a UV laser for the generation of photon pairs in the visible spectrum, while the use of FWM allows the generation of pairs with frequencies close to any excitation laser. This peculiarity means that using FWM can have more versatility in the choice of optics for entangled photon pair generation.
@@ Acknowledgments
This work was supported by IDOR/Pioneer Science Initiative (www.pioneerscience.org), CNPq (INCT-IQ 465469/2014-0, 302872/2019-1, 421469/2023-4, 307619/2023-0), FAPEMIG (APQ-01860-22, APQ-04852-23, RED0008123) and FAPERJ (CNE E-26/200.307/2023).

### v3

@@ IV Polarization entanglement in the SaS pairs
With the values of $A^{\eta}_{ijkl}$ in hand, one can predict what is the polarization entanglement in the SaS scattered state $|\Psi_{SaS}\rangle(\theta)$ of Eq. (5) for given values of $\omega_{S}$ and $\omega_{aS}$. We use, as an entanglement measure, the entropy of entanglement $E=-\text{Tr}_{i}(\rho_{i}\text{log}_{2}\rho_{i})$ of either subsystem $i=\{S,aS\}$ with reduced state $\rho_{i}$ [17], which, since our global state is pure, is zero for separable states and unity for maximally entangled states.
@@
Figure 3: Entanglement map for the SaS states in (a) the present experiment and (b) a generic centrosimetric cubic system. In (a), the laser bandwidth $W$ is fixed at $W=264\text{ cm}^{-1}$, the value of our experiment, as localized in (b) by a horizontal line. The black vertical line is the resonant Raman shift $\omega_{ph}/2\pi$, and the red regions mark the entanglement maxima ($E>0.999$). As tensor components in Eq. (6) we use the values from Table 1. The star localizes the experiment of Ref. [12]. In blue shades, we plot contour curves for the Gisin parameter $F$. The hatched regions, within $2W$, indicate a high production of independent SaS photon pairs.
@@
In Fig. 3 we show a contour plot of $E(\omega_{S},\omega_{aS};\theta,W)$, under the condition that $\omega_{aS}=2\omega_{c}-\omega_{S}$, that is, a symmetric SaS Raman shift. On the horizontal axis we vary the Raman shift, going through the resonance in diamond at $\omega_{ph}/2\pi=1332$ cm-1, identified by a black vertical line, and on the vertical axis we plot (a) the angle $\theta$ between the laser linear polarization and the crystallographic axis, going from $\theta=0^{\circ}$ to $45^{\circ}$ (due to the symmetry of the crystal, all other $\theta$ values will be related to this range), and (b) the laser bandwidth $W$, going from zero to $W=190\gamma$, and where the value for our experiment $W=24\gamma$ ($W=264\text{ cm}^{-1}$ or FWHM 70 cm-1) is indicated by a black horizontal line. The value of $\gamma=11\text{ cm}^{-1}$ is obtained experimentally using a CW laser. The thick red regions indicate the parameters where $E\rightarrow 1$, showing maximum entanglement.
@@
In Fig. 3 there is a hatched region indicating where the uncorrelated SaS pair production is high (low $g^{(2)}$ region in Fig. 2 (c,d)), corresponding to $2W$, around $1.2\times$ FWHM. In this region, the scattered state is not properly represented by Eq. (II), but it needs to be complemented by other SaS scattering events, which involves the scattering of real phonons, that happens within a laser bandwidth around the Raman resonance [18]. Outside this region, though, state (II) is a good representation of the scattered SaS state, and our entanglement measure is representative.
@@
To complete the analysis, we also draw level curves for the violation of Bell-type CHSH inequalities as a function of the light-matter parameters. These curves correspond to the so-called Gisin parameter $F$ [19, 20] which, for pure states, reads $F=2(1+2\mathcal{P})^{1/2}$, where $\mathcal{P}=1-\text{Tr}_{i}(\rho_{i})^{2}$ is the linear entropy of either subsystem $i$. For separable states $F=2$, which means that no separable state can go above the classical upper bound in the CHSH inequality, equal to 2. It reaches its maximum at $F=2\sqrt{2}\approx 2.83$ for maximally entangled states. We plot the $F$ parameter in Fig. 3 as blue shaded contour curves, and its smallest values within a white contour. Note that the maximal violation coincides with the maximally entangled states near 1200 cm-1 and 1400 cm-1. Since Bell analysis is state dependent, and the generated $|\Psi_{SaS}\rangle$ state depends on the values of the light-matter parameters, the appropriate Bell angles to reach the maximum violation change for different regions of the map. As an example, the violation in the maximally entangled states near 1200 cm-1 is obtained for linear polarization angles $(0,\pi/4)$ for one of the photons and $(\pi/8,3\pi/8)$ for the other, where the angles are in relation to $V$ polarization. Note that the quality of the entanglement, characterized by both the Gisin parameter $F$ and the entanglement entropy $E$, are independent of the rate of pair generation, so no matter how many SaS photon pairs are generated, our analysis yields how much entanglement will be available with each pair.
@@
Finally, we can localize the result of Ref. [12] in Fig. 3, represented by a white star. In the paper, a state close to a Bell state at a symmetric Raman shift of 900 cm-1 with $\theta=0^{\circ}$ was found, with amplitude ratio of $|HH\rangle$ to $|VV\rangle$ at $\sqrt{0.28/0.72}\approx 0.62$. Calculating the ratio with our state (5) yields 0.49, and the discrepancy comes from our anti-Stokes filter cutting some of the Raman contribution in HH at 900 cm-1 (left edge of Fig. 2 (a)), which does not happen in Ref. [12]. This region can be seen in the bottom left of Fig. 3 (a) and (b), where $E\approx 0.7$ and $F\approx 2.5$.
V Conclusion
@@
This work presents a fully quantum theory that describes the broadband scattered state in the SaS process, including photon pair polarization. It incorporates experimental conditions like the laser bandwidth and crystal orientation, to allow for the calculation of any sought physical quantity (e.g. $g^{(2)}(0)$ and entanglement). The efficiency of the process can be controlled by how close the system is to the resonance with a Raman active phonon, and it is a matter of state engineering to choose the appropriate laser–crystal angle and SaS frequency to obtain a good balance between a good amount of entanglement and a sufficient pair production rate. This theory has been elusive to find despite a decade-long effort [2, 10, 15, 21, 22, 12]. The result is explored here for diamond, but it should be general for other centrosymmetric media, including silicon, whose electronic band gap (1.12 eV) is one order of magnitude higher than the phonon energy (<0.1<0.1 eV), respecting the assumptions of our model. This should considerably bolster the engineering of SaS photon pair generation. We note that related results were independently obtained in Ref. [23].
@@
For nonzero $\chi^{(2)}$ materials, spontaneous parametric down-conversion (SPDC) enables broadband pair generation as well, as has been recently demonstrated [25]. However, in SPDC the generation of a polarization entangled photon pair is due to the decay of a single highly energetic photon [5], demanding a UV laser for the generation of photon pairs in the visible spectrum, while the use of four-wave mixing (FWM) allows the generation of pairs with frequencies close to any excitation laser. This peculiarity means that using FWM can have more versatility in the choice of optics for entangled photon pair generation.
@@ Acknowledgments
We acknowledge Professor C. Galland and his team for insightful discussions. This work was supported by IDOR/Pioneer Science Initiative (www.pioneerscience.org), Conselho Nacional de Desenvolvimento Científico (INCT-IQ 465469/2014-0, 302872/2019-1, 421469/2023-4, 307619/2023-0), Fundação de Amparo à Pesquisa do Estado de Minas Gerais (APQ-01860-22, APQ-04852-23, RED0008123) and Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (CNE E-26/200.307/2023, E-26/210.069/2020).
Appendix A Analysis of the experimental data
In this Appendix we provide the details of the experimental data and its analysis. The experimental data is composed of a collection of temporal difference histograms associated with frequency filtered Stokes detections. The Stokes signal, filtered with the monochromator, is scanned around the diamond’s first order Raman peak (1332 cm-1) in a spectral region ranged from 895 to 1750 cm-1. For each monochromator position, one histogram of temporal difference is obtained in $300$ seconds of acquisition time. This scanning procedure is done for two positions of the sample ($\theta=0^{\circ}$ and $45^{\circ}$) and for each sample position the spectrum is obtained with the polarization of the photons incident on the APDs being selected in two ways: a spectrum with Stokes and anti-Stokes photons with vertical polarization (VV) and a spectrum with both photons with horizontal polarization (HH).
From the histograms two kinds of experimental data are obtained, the intensity of correlated pairs $I^{corr}_{SaS}(\Delta\tau=0)$, and the second-order normalized correlation function $g^{(2)}(\tau=0)$, respectively given by
$I^{corr}_{SaS}(\Delta\tau=0)$ $=$ $I_{SaS}(\Delta\tau=0)-\bar{I}_{SaS}(\Delta\tau\neq 0),$ (8)
$g^{(2)}(\Delta\tau=0)$ $=$ $\frac{I_{SaS}(\Delta\tau=0)}{\bar{I}_{SaS}(\Delta\tau\neq 0)},$ (9)
where $I_{SaS}(\Delta\tau=0)$ is the total coincident counts of SaS photons pairs in $\Delta\tau=0$ and $\bar{I}_{SaS}(\Delta\tau\neq 0)$ is the average of uncorrelated S,aS photons pairs in $\Delta\tau\neq 0$. These data are obtained from each experimental point in the spectral range and corrected for the detection efficiency of the system (see Appendix B).
The count of correlated Stokes–anti-Stokes (SaS) photon pairs was experimentally obtained using the setup shown in Fig. 1. The PicoHarp 300 correlator provides a histogram of the detection rate of photon pairs as a function of time. An example is shown in Fig. 4. The peaks of coincidences are separated by 13 nanoseconds due to the time rate of the pulsed laser. With the 0.128 picosecond resolution of this correlator, and considering the time jitter of both APDs (Avalanche Photodiodes) at 350 picoseconds, we have three time bins of points before and after the peak value. This is type of the histogram integration we use to improve the statistics of the photon counts. All data values were obtained from a 300 seconds measurement.
Figure 4: (a) Typical histogram for counts of SaS photon pairs created by the PicoHarp 300 is shown, with the utilized values for integrating the peaks highlighted in yellow. (b) Peak for $\Delta\tau=0$ zoomed in to clarifying the yellow points.
For each data histogram, we determine the correlated rate of SaS photon pairs using the peak at $\Delta\tau=0$, and the uncorrelated rate using the average of the peaks at $\Delta\tau\neq 0$. For intensities of correlated SaS photon pairs, $I_{SaS}(\Delta\tau=0)$, we can create a spectral graph of the obtained values, as shown in Fig. 5 (a) and (b) for crystallographic orientations of 0° and 45°, respectively, with VV (blue dots) and HH (red dots) polarization. The experimental values related to the average of the uncorrelated SaS photon pairs, $\bar{I}_{SaS}(\Delta\tau\neq 0)$, can be demonstrated in Fig. 5 (c) for the sample oriented at 0° with VV and HH polarization (following the same color scheme as the other figures) and in Fig. 5 (d) for the diamond oriented at 45°. We observed that in all data the values of $I_{SaS}(\Delta\tau=0)$ show a peak in the resonance region, which can be attributed to an increase in uncorrelated SaS photon pairs in this region, as shown in Fig. 5 (c,d).
Figure 5: Experimental graphs without efficiency corrections for the intensity of correlation pairs, $I_{SaS}(\Delta\tau=0)$, per second for (a) 0 and (c) 45 crystallographic orientations using VV and HH polarizations. The other graphs show the uncorrelated SaS photon pairs, $\bar{I}_{SaS}(\Delta\tau\neq 0)$, per second for VV polarization (blue dots) and HH polarization (red dots), for the crystallographic orientation of (c) 0 and (d) 45.
By subtracting $\bar{I}_{SaS}(\Delta\tau\neq 0)$ from $I_{SaS}(\Delta\tau=0)$, we obtain the intensity of correlated SaS photons, $I^{corr}_{SaS}(\Delta\tau=0)$. In Fig. 6, we show the curves for 0° (Fig. 6 (a)) and 45° (Fig. 6 (b)) for VV and HH polarizations. The values here differ from those shown in Fig. 2 due to efficiency corrections made to the data, but the behavior of the curves remains the same.
Figure 6: Graphs for the intensities of correlated SaS photon pairs, $I^{corr}_{SaS}(\Delta\tau=0)$, without efficiency corrections using VV and HH polarizations for (a) 0 and (b) 45 angular positions of diamond.
The error bars of the uncorrelated counts (Fig. 5 (c) and 5 (d)) were estimated in two steps, by noting that the curves consist of a central Gaussian peak plus a constant background. The error of the constant background is two times the standard deviation of the counts around the mean. On the other hand, the error of the Gaussian is obtained by fitting a Gaussian curve to the data with the Python SciPy library, and then the error bars correspond to the uncertainty in the fit, provided by the library.
The error bars of the correlated counts in Fig. 5 (a), red curve, and Fig. 5 (b), blue curve, were obtained from the curves without the Raman contribution, namely VV at 0 and HH at 45, just like we did with the uncorrelated counts, that is, separating the flat background and the Gaussian peak contribution. Then we assume that the error of the HH or VV counts are of the same order at 0 and at 45, that is, we use the same error bars of VV at 0 in the curve of Fig. 5 (a) in the VV(45) data in Fig. 5 (b). Similarly, we use the same error bars of HH at 45 in the curve of Fig. 5 (b) in the HH(0) data in Fig. 5 (a).
The error bars for the correlated counts $I_{SaS}^{corr}(\Delta\tau=0)=I_{SaS}(\Delta\tau=0)-\bar{I}_{SaS}(\Delta\tau\neq 0)$ in Fig. 2 were obtained by error propagation of independent variables from both $\Delta\tau=0$ counts $I_{SaS}(\Delta\tau=0)$ of Fig. 5 (a,b), and $\Delta\tau\neq 0$ counts $\bar{I}_{SaS}(\Delta\tau\neq 0)$ counts of Fig. 5 (c,d), according to
$\delta I_{SaS}^{corr}(\Delta\tau=0)=\sqrt{[\delta I_{SaS}(\Delta\tau=0)]^{2}+[\delta\bar{I}_{SaS}(\Delta\tau\neq 0)]^{2}},$ (10)
where $\delta x$ indicates the uncertainty of quantity $x$. In Fig. 6, the errors are the same as the ones in Fig. 2 (a,b), but without efficiency correction.
The error bars of $g^{(2)}(0)$ in Fig. 2 (c,d) were obtained by standard error propagation of independent variables from the $\Delta\tau=0$ counts $I_{SaS}(\Delta\tau=0)$ of Fig. 5 (a,b), and $\Delta\tau\neq 0$ counts $\bar{I}_{SaS}(\Delta\tau\neq 0)$ counts of Fig. 5 (c,d), with $g^{(2)}(0)=I_{SaS}(\Delta\tau=0)/\bar{I}_{SaS}(\Delta\tau\neq 0)$, yielding
$\delta g^{(2)}(0)=\sqrt{\left[\frac{1}{\bar{I}_{SaS}(\Delta\tau\neq 0)}\right]^{2}[\delta I_{SaS}(\Delta\tau=0)]^{2}+\left[\frac{I_{SaS}(\Delta\tau=0)}{(\bar{I}_{SaS}(\Delta\tau\neq 0))^{2}}\right]^{2}[\delta\bar{I}_{SaS}(\Delta\tau\neq 0)]^{2}}.$ (11)
Appendix B Efficiency Correction
Optical measurements are affected by the efficiency of the optical elements used in the experimental setup. Therefore, to ensure that the measured values are closer to an ideal measurement—i.e., without photon losses along the optical path and during detection—a correction was applied to the values in Fig. 6. This correction was based on the product of all the efficiencies of the optical elements used in the experiment, with the final value as a function of the Raman shift exemplified in Fig. 7 (a). All efficiency values were obtained from the manufacturer’s website.
In addition to this correction, the fact that fluctuations in laser power can occur was taken into account. Therefore, the laser power was measured every ten measurements. These values are specified in Fig. 7 (b) for the data related to the crystallographic orientation of 0° for VV and HH, and in Fig. 7 (c) for the 45° orientation for both polarizations.
From these values, the correction
$\mbox{Real Value}=\frac{\mbox{Measured Value}}{\mbox{Efficency}\times\mbox{Laser Power}^{2}}$ (12)
is obtained for each measured value. Note that in Eq. (12) the laser power is squared due to the quadratic nature of SaS photon pair formation with the laser power.
Figure 7: (a) Efficiency of the total optical elements used as a function of Raman Shift in the experimental setup of Fig. 1 (all values were obtained from the manufacturer’s website). Moreover, this figure shows the spectrum of the laser power in miliwatt for (b) $\theta=0^{\circ}$ and (c) $\theta=45^{\circ}$ for VV and HH polarizations.
Appendix C Obtaining the material tensor components
The values of the material tensor components in Table 1 are calculated using the plot fitting parameters $Y^{E}_{HH}(0^{\circ})=\langle I_{SaS}^{corr}(\text{VV}(0^{\circ}))\rangle^{1/2}(0.68-i0.12)$, $Y^{E}_{VV}(45^{\circ})=\langle I_{SaS}^{corr}(\text{VV}(0^{\circ}))\rangle^{1/2}(1.61-i0.55)$, and $Y^{R}_{HH}(0^{\circ})=Y^{R}_{VV}(45^{\circ})=51450\ [\text{Hz/W}^{2}]^{1/2}$, and Eq. (6).
According to group theory analysis, the VV(0) configuration does not exhibit a Raman contribution ($A^{R}_{xxxx}=0$ in Eq. (6)) and it only involves non-resonant electronic transitions. For this reason, the blue curve shown in Fig. 2 (a) of the article corresponding to this configuration is flat. Additionally, Fig. 2 (a) shows a dark blue solid line, obtained by averaging the intensity values $I_{SaS}^{corr}$ for this experimental configuration, and the standard deviation of the experimental points around this mean is at 13%. With the mean and standard deviation of VV(0) counts, times the monochromator slit spectral width $\Delta\omega/2\pi=11\text{ cm}^{-1}$, at $\langle I_{SaS}^{corr}(\text{VV}(0^{\circ}))\rangle=(27.5\pm 3.5)\times 10^{3}\text{ Hz/W}^{2}$, we take its square root and use $\langle I_{SaS}^{corr}(\text{VV}(0^{\circ}))\rangle^{1/2}=MY^{E}_{VV}(0^{\circ})=(166\pm 11)\ [\text{Hz/W}^{2}]^{1/2}$ as our reference value, where $M=0.551$ is a constant accounting for the integration of $f^{E}(\omega_{S},\omega_{aS})$ over the filter frequencies (see Eq. (5) and (7)). The uncertainty is obtained by error propagation. Since $Y^{E}_{VV}(0^{\circ})=C\mathcal{E}_{0}^{2}A^{E}_{xxxx}$, we have a reference value for $A^{E}_{xxxx}$.
We can now look at the HH(0) data in Fig. 2 (a), represented by the red points, which contains p-SaS, as can be seen from the characteristic resonance signature. In this configuration we have $A^{R}_{xyyx}\neq 0$ in Eq. (6), contributing with a Raman scattering in the $|VV\rangle$ polarization, and which we will calculate. To fit the data with Eqs. (II)–(6), we used $Y^{E}_{HH}(0^{\circ})=MY^{E}_{VV}(0^{\circ})(0.68-i0.12)$ and $Y^{R}_{HH}(0^{\circ})=51450\ [\text{Hz/W}^{2}]^{1/2}$. Since $Y^{\eta}_{HH}(0^{\circ})=C\mathcal{E}_{0}^{2}A^{\eta}_{xyyx}$, these numbers lead to
$A^{E}_{xyyx}/A^{E}_{xxxx}=(0.37\pm 0.04)+i(-0.07\pm 0.01)$
and
$A^{R}_{xyyx}/A^{E}_{xxxx}=(171\pm 12).$
Note that $|A^{E}_{xyyx}|/|A^{E}_{xxxx}|$ is close to the $1/3$ expected from group theory [13] and we have a 3% absorption in the electronic degrees of freedom. Also, the electronic-to-Raman ratio $A^{E}_{xyyx}/A^{R}_{xyyx}\approx 0.002$ should be comparable to the susceptibilities ratio near the resonance since, by $\chi^{(3)R}_{ijkl}=A^{R}_{ijkl}\gamma/(\omega_{ph}-\omega_{L}+\omega_{S}+i\gamma/2)$, at resonance $|\chi^{(3)R}_{ijkl}|=2|A^{R}_{ijkl}|$.
The HH(45) configuration also exhibits only e-SaS, as seen in the red data in Fig 2 (b) of the main article. The mean and standard deviation of the counts gives $(1.60\pm 0.28)\times 10^{3}\text{ Hz/W}^{2}$, which means a 17% fluctuation. Taking the square root of the counts give us $MY^{E}_{HH}(45^{\circ})=(40\pm 4)\ [\text{Hz/W}^{2}]^{1/2}$. Using $MY^{E}_{HH}(45^{\circ})=MC\mathcal{E}_{0}^{2}[A^{E}_{xxxx}+A^{E}_{xyyx}-(A^{E}_{xxyy}+A^{E}_{xyxy})]/2$ and the values obtained with the VV(0) and HH(0) data, we find
$(A^{E}_{xxyy}+A^{E}_{xyxy})/A^{E}_{xxxx}=(0.89\pm 0.10)+i(-0.07\pm 0.01).$
This value does not coincide with the $2/3$ expected, suggesting that there is some unaccounted effect, which Ref. [13] attributes to some two-photon absorption processes. Furthermore, for the p-SaS contribution in this configuration to be zero, that is, $Y^{R}_{HH}(45^{\circ})=0$, we need $(A^{R}_{xxyy}+A^{R}_{xyxy})=A^{R}_{xyyx}$ (see Eq. (6b)).
The VV(45) data is the last configuration left, and it is shown in Fig. 2 (b) of the main article in blue. When finding the theoretical curve that best fits the data, we fix the Raman factor to be the same as in the HH(0) configuration, $Y^{R}_{VV}(45^{\circ})=Y^{R}_{HH}(0^{\circ})=51450\ [\text{Hz/W}^{2}]^{1/2}$, which should be the case since $Y^{R}_{VV}(45^{\circ})=C\mathcal{E}_{0}^{2}[A^{E}_{xyyx}+(A^{E}_{xxyy}+A^{E}_{xyxy})]/2=Y^{R}_{HH}(0^{\circ})$. The best fitting curve, shown in Fig. 2 (b) of the article, is is obtained by choosing $Y^{E}_{VV}(45^{\circ})=MY^{E}_{VV}(0^{\circ})(1.61-i0.55)$, which gives $MY^{E}_{VV}(45^{\circ})=[(147\pm 10)+i(-50\pm 4)]\ [\text{Hz/W}^{2}]^{1/2}$.
We could, however, have taken advantage of the tensorial relations and calculate via the expression $Y^{E}_{VV}(45^{\circ})=C\mathcal{E}_{0}^{2}[A^{E}_{xxxx}+A^{E}_{xyyx}+(A^{E}_{xxyy}+A^{E}_{xyxy})]/2$, whose components we have already calculated, yielding $MY^{E}_{VV}(45^{\circ})=[(188\pm 15)+i(-12\pm 4)]\ [\text{Hz/W}^{2}]^{1/2}$. We attribute the discrepancy to the difficulty in extracting the correct $(A^{E}_{xxyy}+A^{E}_{xyxy})/A^{E}_{xxxx}$ ratio from the low HH(45) counts. Since the counts are low, the uncertainty is most probably underestimated, and a lower $(A^{E}_{xxyy}+A^{E}_{xyxy})/A^{E}_{xxxx}$ would better fit the VV(45) counts.